OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point 2(x^2+y^2)^2=25(x^2-y^2) and the point (3,1) I've been working on this problem for a long time now and I can't figure out where I'm going wrong.

OpenStudy (anonymous):

find f'(x) first

OpenStudy (anonymous):

isolate y

OpenStudy (anonymous):

plug in x to find the slope of the tangent line

OpenStudy (anonymous):

than write ur equation in the format of (y-y1) = m(x-x1)

OpenStudy (anonymous):

I started by doing the power rule and getting (4x^2+4y^2)[2x+2y dy/dx]=25(2x-2y dy/dx) and then distributed,isolated dy/dx, etc, but my answer is wrong because they are getting a negative answer and from the derivative I get I wouldn't get a negative answer.

OpenStudy (anonymous):

u sure u know the product rule?

OpenStudy (anonymous):

g*f'+f*g'

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

u know ur power and chain rules?

OpenStudy (anonymous):

Yes I set it up correctly that I can see. It is in the few comments above this

OpenStudy (anonymous):

well can i see ur isloated f(x)?

OpenStudy (anonymous):

isolated*

OpenStudy (anonymous):

I know it is wrong because the answer is negative and my dy/dx isn't going to give a negative answer but I got x/(8x^3+8x^2y+8y^2x+8y^3+y)

OpenStudy (anonymous):

all of that equals y? hly smokes...

OpenStudy (anonymous):

I mean that's what I got but I don't know. I don't see where I went wrong, I just know it isn't correct.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

ur y isn't isolated...

OpenStudy (anonymous):

you are isolating dy/dx (the derivative of y with respect to x), not the y variable.

OpenStudy (anonymous):

or y', it means the same thing

OpenStudy (anonymous):

mhmm ok

OpenStudy (anonymous):

gtg soon so can't solve the whole thing...

OpenStudy (loser66):

back to your previous post to see